The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.
Answer by checkley71(8403)

PERIMETER OF A RECTANGLE IS 2*L+2*W THEREFORE 2(2X+2)+2*X=34 OR
4X+4+2X=34 OR 6X=30 OR X=30/6 OR X=5 PROOF
4(2*5+2)+2*5=34 OR 4(10+2)+10=34 OR 4*12+10=34 OR 24+10=34 OR 34=34